# Griesmer and Optimal Linear Codes from the Affine Solomon-Stiffler Construction

### v2

@@ 3 Optimal few-weight affine Solomon-Stiffler codes
Let us consider the case $k=2m$, $u_{1}=u_{2}=m$. Then we construct a linear $[q^{2m}-2q^{m}+1,2m,(q-1)(q^{2m-1}-2q^{m-1})]_{q}$ code. From Theorem 2, the Griesmer defect is one and this is an optimal code. These codes have the same parameters as optimal codes constructed by Shi, Guan and Solé in [27, Theorem 6.1]. The optimal codes in [27, Theorem 6.2], are just a repetition copy of the above code. Their $q$ needs to be an odd prime satisfying $q\equiv 3$ $od$ $4$, and $m$ needs to be odd. It should be indicated that our codes are more general. Moreover optimal codes in [27] were constructed via Gray images of two-weight codes over a finite ring.
@@ 00 $1$
Notice that the weight distribution of this infinite family of optimal two-weight $q$-ary linear codes is the same as the weight distribution Table II of optimal two-weight codes constructed in [27, Theorem 6.1]. Their construction via Gray images of two-weight codes over the finite ring ${\bf F}_{p}+u{\bf F}_{p}$, seems unnecessary, since their strong conditions on $q$ and $m$ are not essential. When $m=2r$, codes constructed in [16, Proposition 5] are these codes in Theorem 3.
@@
Table 3: Weight distribution of optimal two-weight codes
@@
Table 4: Weight distribution of optimal two-weight codes
@@
Table 5: Weight distribution of optimal two-weight codes
@@ 4 Modified affine Solomon-Stiffler codes
In affine Solomon-Stiffler codes constructed in Theorem 1, if a vector ${\bf v}$ is a column vector of the generator matrix $G$, then $\lambda{\bf v}$ is also a column vector of $G$. If we only pick up these columns of the form $\lambda{\bf v}$, where $\lambda$ takes over all elements of a subgroup of the order $e$ in the multiplicative group ${\bf F}_{q}^{*}$, we have the following result.
@@
The interesting point here is that Griesmer defects of these modified affine Solomon-Stiffler codes decrease. It is easy to see that when $e=1$ and $u_{1}<u_{2}<\cdots<u_{h}$. Our modified affine Solomon-Stiffler codes are $q$-ary projective Solomon-Stiffler codes invented in [30].
@@
Then the repetiton two-copy of this code is a linear Griesmer $[\frac{2e(q^{2m}-q^{m}}{q-1},2m,2e(q^{2m-1}-q^{m-1})]_{q}$ code, with the weight distribution
@@
Notice that this is the same as the Griesmer codes constructed in [19, Theorem 3.5]. The weight distribution is the same as that in [19, Theorem 3.3]. The construction in [19], as that in [27], is very technique, and some restrictions on $e$ and $m$ were needed.
@@
In the following table, we list some non-Solomon-Stiffler codes constructed from Theorem 6. Most of them are optimal, see [9].
@@ 00 $1$
Table 8: Weight distribution of optimal two-weight codes
@@ 6 Modified affine Solomon-Stiffler codes when $u_{1}=u_{2}\cdots=u_{h}=1$, $h$ large
Proof. From the proof of Theorem 1, the conclusion is equivalent to find a nonzero vector vector ${\bf h}$ such that the inner products of ${\bf h}$ with these $h$ lines are nonzero. This is equivalent to the existence of a projective linear $[h,k]_{q}$ code with the maximum weight $h$.
@@ 7 Distance-optimal codes as codimension one or two subcodes in Griesmer codes
Theorem 9 Let ${\bf C}$ be a binary linear Griesmer $[n,k,2d]_{2}$ code satisfying $k\geq 3$ and $d\leq 2^{k-2}$. Then any dimension $k-1$ subcode of ${\bf C}$ has minimum weight $2d$ and is distance-optimal. The Griesmer defect of this shortening code is $1$.
@@
Distance-optimal codes constructed in [26, Theorem 2] and [25, Proposition 5.3, 5)] have the same parameters as above codes. Their constructions are complicated and technique. Actually, more general such distance-optimal binary codes are constructed in the following result. By the way, optimal binary three-weight $[3\cdot 2^{3m-1},3m,3\cdot 2^{3m-2}]_{2}$ codes constructed in [29] have the same parameters as repetition codes of the 1st order RM $[2^{3m-1},3m,2^{3m-2}]_{2}$ codes.
@@
Proof. If the codimension two subcode is not distance-optimal. Then there exists a linear $[n,k-2,4d+1]_{2}$ code. From the Griesmer bound,
@@ 9 Conclusion
In this paper, we propose a general construction of Griesmer and optimal linear codes as an affine version of Solomon-Stiffler codes. Modified affine Solomon-Stiffler codes include these projective Solomon-Stiffler codes as special cases. Griesmer or optimal codes constructed in [27, 19] have the same parameters and weight distributions as very special case codes in our construction. Our Griesmer and optimal codes are less restrictive than Griesmer and optimal trace codes constructed in [16]. Optimal binary codes constructed in a recent paper [25] and another paper [26] and codimension one subcodes of special binary Solomon-Stiffler codes have the same parameters. Distance-optimal codes as punctured code of Griesmer codes were also discussed. We reconstructed many optimal codes in [9] from our modified affine Solomon-Stiffler construction. Projective linear almost optimal codes similar to codes in [30] were also constructed.

### v3

@@ 3 Optimal few-weight affine Solomon-Stiffler codes
Let us consider the case $k=2m$, $u_{1}=u_{2}=m$. Then we construct a linear $[q^{2m}-2q^{m}+1,2m,(q-1)(q^{2m-1}-2q^{m-1})]_{q}$ code. From Theorem 2, the Griesmer defect is one and this is an optimal code. These codes have the same parameters as optimal codes constructed by Shi, Guan and Solé in [36, Theorem 6.1]. The optimal codes in [36, Theorem 6.2], are just a repetition copy of the above code. Their $q$ needs to be an odd prime satisfying $q\equiv 3$ $od$ $4$, and $m$ needs to be odd. It should be indicated that our codes are more general. Moreover optimal codes in [36] were constructed via Gray images of two-weight codes over a finite ring.
@@ 00 $1$
Notice that the weight distribution of this infinite family of optimal two-weight $q$-ary linear codes is the same as the weight distribution Table II of optimal two-weight codes constructed in [36, Theorem 6.1]. Their construction via Gray images of two-weight codes over the finite ring ${\bf F}_{p}+u{\bf F}_{p}$, seems unnecessary, since their strong conditions on $q$ and $m$ are not essential. When $m=2r$, codes constructed in [20, Proposition 5] are these codes in Theorem 3.
@@
Table 3: Weight distribution of optimal three-weight codes
@@
Table 4: Weight distribution of optimal three-weight codes
@@
Table 5: Weight distribution of optimal five-weight codes
@@ 4 Modified affine Solomon-Stiffler codes
In affine Solomon-Stiffler codes constructed in Theorem 1, if a vector ${\bf v}$ is a column vector of the generator matrix $G$, then $\lambda{\bf v}$, where $\lambda\in{\bf F}_{q}^{*}$, is also a column vector of $G$. If we only pick up these columns of the form $\lambda{\bf v}$, where $\lambda$ takes over all elements of a subgroup of the order $e$ in the multiplicative group ${\bf F}_{q}^{*}$, we have the following result.
@@
The interesting point here is that Griesmer defects of these modified affine Solomon-Stiffler codes decrease. It is easy to see that when $e=1$ and $u_{1}<u_{2}<\cdots<u_{h}$, modified affine Solomon-Stiffler codes are $q$-ary projective Solomon-Stiffler codes invented in [39].
@@
Then the repetition two-copy of this code is a linear Griesmer $[\frac{2e(q^{2m}-q^{m})}{q-1},2m,2e(q^{2m-1}-q^{m-1})]_{q}$ code, with the weight distribution
@@
Notice that this is the same as the Griesmer codes constructed in [23, Theorem 3.5]. The weight distribution is the same as that in [23, Theorem 3.3]. The construction in [23], as that in [36], is technical, and some restrictions on $e$ and $m$ were needed.
@@
In the following table, we list some non-Solomon-Stiffler codes constructed from Theorem 6. Most of them are optimal, see [12]. These codes have few nonzero weights, and their exact weight distributions can be determined explicitly.
@@ 00 $1$
Table 8: Weight distribution of optimal three-weight codes
@@ 6 Modified affine Solomon-Stiffler codes when $u_{1}=u_{2}\cdots=u_{h}=1$, $h$ large
Proof. From the proof of Theorem 1, the conclusion is equivalent to finding a nonzero vector vector ${\bf h}$ such that the inner products of ${\bf h}$ with these $h$ lines are nonzero. This is equivalent to the existence of a projective linear $[h,k]_{q}$ code with the maximum weight $h$.
@@ 7 Distance-optimal codes as codimension one or two subcodes in Griesmer codes
Theorem 9 Let ${\bf C}$ be a binary linear Griesmer $[n,k,2d]_{2}$ code satisfying $k\geq 3$ and $d\leq 2^{k-2}$. Then any dimension $k-1$ subcode of ${\bf C}$ has minimum weight $2d$ and is distance-optimal. The Griesmer defect of this subcode is $1$.
@@
Distance-optimal codes constructed in [35, Theorem 2] and [30, Proposition 5.3, 5)] have the same parameters as above codes. Their constructions are complicated and technique. Actually, more general such distance-optimal binary codes are constructed in the following result. By the way, optimal binary three-weight $[3\cdot 2^{3m-1},3m,3\cdot 2^{3m-2}]_{2}$ codes constructed in [38] have the same parameters as repetition codes of the 1st order RM $[2^{3m-1},3m,2^{3m-2}]_{2}$ codes.
@@
Proof. If the codimension two subcode is not distance-optimal, then there exists a linear $[n,k-2,4d+1]_{2}$ code. From the Griesmer bound,
@@ 9 Conclusion
In this paper, we propose a general construction of Griesmer and optimal linear codes as an affine version of Solomon-Stiffler codes. Modified affine Solomon-Stiffler codes including projective Solomon-Stiffler codes as special cases, were given. Griesmer or optimal codes constructed in [36, 23] have the same parameters and weight distributions as very special case codes in our construction. Optimal binary codes constructed in a recent paper [30] and another paper [35] and codimension one subcodes of special binary Solomon-Stiffler codes have the same parameters. Distance-optimal codes as punctured or shortened codes of Griesmer codes were also discussed. We reconstructed many optimal codes in [12] from our modified affine Solomon-Stiffler codes. Projective linear almost optimal codes similar to codes in [39] were also constructed.